THE PREDICTED INFRARED SPECTRUM OF THE HYPERMETALLIC MOLECULE CaOCa IN ITS LOWEST TWO ELECTRONIC STATES X̃ ^{1}Σ_{g}^{+} AND ã ^{3}Σ_{u}^{+}

B. OSTOJIĆ, Institute of Chemistry, Technology and Metallurgy, University of Belgrade, Studentski trg 14-16, 11 000 Belgrade, Serbia; P. R. BUNKER, P. SCHWERDTFEGER, Centre for Theoretical Chemistry and Physics (CTCP), The New Zealand Institute for Advanced Study(NZIAS), Massey University Auckland, Private Bag 102904, North Shore City, 0745 Auckland, New Zealand; ARTUR GERTYCH, \underline{PER JENSEN}, FB C – Physikalische und Theoretische Chemie, Bergische Universität, D-42097 Wuppertal, Germany.

This study of CaOCa is the third in a series by us on Group 2 alkaline-earth M_{2}O hypermetallic oxides. As with our previous calculations in the series, on BeOBe and MgOMg, the ab initio calculations we report here show that CaOCa has a linear ^{1}Σ_{g}^{+} ground electronic state and a very low lying linear ã ^{3}Σ_{u}^{+} first excited triplet electronic state. For CaOCa we determine that the singlet-triplet splitting T_{e}(ã) = 386 cm^{−1}. We calculate the three-dimensional potential energy surface, and the electric dipole moment surfaces, of each of the two states using a multireference configuration interaction (MRCISD) approach in combination with internally contracted multireference perturbation theory (RS2C) based on fullvalence complete active space self-consistent field (FV-CASSCF) wavefunctions with a cc-pwCVQZ-DK basis set for Ca and a cc-pCVQZ basis set for O. We simulate the infrared absorption spectra of ^{40}Ca^{16}O^{40}Ca in each of these electronic states in order to aid in its eventual spectroscopic characterization.